OpenStudy (anonymous):

The probability of a spinner stopping on 5 is . The probability of the spinner stopping on 6 is . What is the probability of the spinner stopping on 5 or 6?

Mehek (mehek14):

could you check your question some parts are missing

OpenStudy (anonymous):

A. B. C. D.

OpenStudy (anonymous):

The probability of a spinner stopping on 5 is 1/9 . The probability of the spinner stopping on 6 is 4/9 . What is the probability of the spinner stopping on 5 or 6

OpenStudy (vera_ewing):

You can add the fractions to get the answer.

Mehek (mehek14):

so just add \(\dfrac{1}{9}+\dfrac{4}{9} = \)

Mehek (mehek14):

just add 1 + 4 leave the denominator as is

OpenStudy (anonymous):

so what is it

OpenStudy (vera_ewing):

\[\frac{ 1 }{ 9 } + \frac{ 4 }{ 9 } =~?\]

Mehek (mehek14):

\(\dfrac{1+4}{9}\)

OpenStudy (vera_ewing):

What is 1+4?

Mehek (mehek14):

add 1 + 4

OpenStudy (anonymous):

5/9

OpenStudy (vera_ewing):

Yes!

Mehek (mehek14):

so that's your answer

OpenStudy (vera_ewing):

Good job :)

OpenStudy (anonymous):

^^